Angles A and B are supplementary angles and the measure of angle A is 2 degrees more than 4 times the measure of angle B. Find the measures of angle A and angle B.
Answer by macston(5194)

The sum of supplementary angles=180 degrees
Let B= one angle and the other A,its supplement=4B+2
A+B=180 degrees Substitute for A.
(4B+2)+B=180 degrees Subtract 2 degrees from each side
5B=178 degrees Divide each side by 5
B=35.6 degrees ANSWER 1: The measure of angle B is 35.6 degrees
A=4B+2=4(35.6)+2=142.4+2=144.4 degrees ANSWER 2: The measure of angle A is 144.4 degrees.
CHECK
The sum of supplementary angles=180 degrees
A+B=180 degrees
144.4 degrees + 35.6 degrees=180 degrees
180 degrees = 180 degrees